A combine costing $32,000.00 has a trade-in value of $5,000.00 after 8 years. Compute the book value after 2 years and the depreciation charge in year 3 for

a. The straight-line method

b. The sum-of-the-years-digits method

c. The simple declining-balance method

d. The complex declining-balance method
Answer #1

a. Straight-line method Depreciation expense per year = (Cost - Salvage Value) / Useful life = (32,000 - 5,000) / 8 = $3,375

b. Sum-of-the-years-digits method SYD = 1+2+3+4+5+6+ 7+8 = 36 Depreciable amount = Cost - Salvage Value = 32,000 - 5,000 = $2

c. Simple declining-balance method SL depreciation rate = (1/ Useful life) * 100 = (1/8) * 100 = 12.5% Simple declining-balan

d. Complex declining-balance method (Double declining-balance method) SL depreciation rate = (1/ Useful life) * 100 = (1/8) *